OpenStudy (cenaida):

What is the y-intercept of the line? Y=(5/3)x-5 A. -5 B. 5

jimthompson5910 (jim_thompson5910):

hint: y = mx+b m = slope b = y-intercept

OpenStudy (cenaida):

So is it 5 or negative 5

jimthompson5910 (jim_thompson5910):

for example, in y = 2x + 3 we see that m = 2 b = 3 so the y-intercept here is 3

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (cenaida):

Yea but i dont know if thats a minus or negative so im not sure if the y intercept is 5 or -5

jimthompson5910 (jim_thompson5910):

well it would be b = 5 if we had y = (5/3)x + 5 but it's actually y = (5/3)x - 5

jimthompson5910 (jim_thompson5910):

so it might help to write y = (5/3)x - 5 as y = (5/3)x + (-5)

OpenStudy (cenaida):

So its -5

jimthompson5910 (jim_thompson5910):

yes